OpenStudy (anonymous):

Please check my answer: I got C. @jim_thompson5910

OpenStudy (anonymous):

looks good to me we can check it with numbers if you like

jimthompson5910 (jim_thompson5910):

C is correct A = L*W 2*A = 2*(L*W) 2*A = L*(2*W) So this shows that if the width (W) and area (A) are doubled to get 2*W and 2*A respectively, then the length must be held the same

OpenStudy (anonymous):

@jim_thompson5910 Thank you , may you check this one. I got C. as well.

jimthompson5910 (jim_thompson5910):

yes the doubling and the halving cancel each other in a way so you're left with the same area

OpenStudy (anonymous):

Thanks :)